OpenStudy (anonymous):

determine if a set of vectors forms a basis for Rn. (Can someone do an example of this please??)

OpenStudy (anonymous):

any bunch of n vector independents

OpenStudy (anonymous):

for R³ could be (1,0,0);(0,1,0);(0,0,1)

OpenStudy (anonymous):

with linear combination of this three vector you can fill the whole R³ space

OpenStudy (anonymous):

A basis is a set of vectors which span the space. As @RaphaelFilgueiras said.

OpenStudy (anonymous):

Given a set of vectors, you would check to see if they span of the vector space in question (i.e. any vector in the given space can be written as a linear combo of the given vectors), and you would check to see if they're linearly independent. Continuing with @RaphaelFilgueiras's example, suppose you're told to show that the set \(\left\{(1,0,0)~,~(0,1,0)~,~(0,0,1)\right\}\) is a basis of \(\mathbb{R}^3\). One of the things you have to do is show \(\text{span}\left\{(1,0,0)~,~(0,1,0)~,~(0,0,1)\right\}=\mathbb{R}^3\). Let \((x,y,z)\in\mathbb{R}^3\) be some vector. Show that \((x,y,z)=c_1(1,0,0)+c_2(0,1,0)+c_3(0,0,1)\) for any \(x,y,z\in\mathbb{R}\). Basically, you have the system of equations \[\begin{cases} c_1=x\\ c_2=y\\ c_3=z \end{cases}\] So, letting \(c_1=x,c_2=y,c_3=z\) means you can indeed write any vector in the space as the linear combination of the given set of vectors. Thus \(\text{span}\left\{(1,0,0)~,~(0,1,0)~,~(0,0,1)\right\}=\mathbb{R}^3\). Lastly, show that the given set of vectors is linearly independent. This basically means you can write the zero vector as a linear combination of the vectors only if the coefficients in front of them are all zero. \[c_1(1,0,0)+c_2(0,1,0)+c_3(0,0,1)=(0,0,0) ~ \iff ~c_1=c_2=c_3=0\] This should be fairly clear for these vectors once you set it up as a system of equations. You'll see that the only possible solutions are \(c_1=c_2=c_3=0\). Thus the set of vectors is linearly independent. And there you have it: \(\left\{(1,0,0)~,~(0,1,0)~,~(0,0,1)\right\}\) is a basis of \(\mathbb{R}^3\).